OpenStudy (epicteatime):

Identify the solutions of 3x2 - 11x - 4 = 0

OpenStudy (epicteatime):

@ganeshie8

OpenStudy (epicteatime):

@jim_thompson5910

OpenStudy (epicteatime):

@myininaya

OpenStudy (anonymous):

try to factorise

OpenStudy (epicteatime):

hmmmmmm

OpenStudy (epicteatime):

-12 x 1 or 12 x -1

OpenStudy (anonymous):

you need -11 so -12 X 1

OpenStudy (epicteatime):

oh ok ;)

OpenStudy (epicteatime):

what next sir?

OpenStudy (anonymous):

3x2 - 12x +x - 4 = 0

OpenStudy (epicteatime):

3x^2-12x+1x-4

OpenStudy (epicteatime):

same thing ;P

OpenStudy (anonymous):

yup

OpenStudy (epicteatime):

whats the answer? do i use the quadratic equation?

OpenStudy (anonymous):

but for eqution right hand side is 0

OpenStudy (anonymous):

3x2 - 12x +x - 4 = 0 so 3x(x-4) +1(x-4) =0 (3x+1)(x-4) =0 here either x-4 =0 or 3x+1 =0 or both hence x=4 or x=-1/3

OpenStudy (epicteatime):

oh it was 1!

OpenStudy (epicteatime):

put 4(x-1) ;P